You have been hired to help estimate the average cattle output in the U.S. for last year. The data is available from a company and a fee is charged for each state sampled. If the range of cattle output by state (in 1000 head) is about 14,000 and they wish to have a 95% confidence interval with a bound of 1,000 for the average, how many states should be sampled? Find solution without using z-score.
Answer #1

Range= 14000

S.d= range/4 =14000/4= 3500

Sample size for single mean

For 95% confidence, z=1.96

E=1000
E = Zα/2 × (σ/√n)
1000 = 1.96 x (3500)/√n
=> n =48